One printer takes 4 hours to complete a job. Another printer can do the same job in 3 hours. When the job runs on both printers, how many hours will it take to complete?
Answer by ptaylor(2198)

Let x=number of hours it takes both printers working together to do the job
One printer prints at the rate of 1/4 of the job per hour
The other printer prints at the rate of 1/3 of the job per hour
Together, they print at the rate of 1/4 + 1/3=3/12 + 4/12=7/12 of the job per hour
So, our equation to solve is:
(7/12)*x=1 (1 job, that is)
7x=12
x=12/7 hr=1 5/7 hr
CK
(1/4)*(12/7)+(1/3)*(12/7)=?1
3/7 + 4/7=1
1=1
Hope this helps--ptaylor